# Selective Self-to-Supervised Fine-Tuning for Generalization in Large Language Models

### v1

@@ 1 Introduction
SFT relies on gold responses for training. We observe two key challenges when performing SFT on LLMs: (1) in many datasets, model-generated responses—though differing from gold responses—are still valid and acceptable, and (2) the distribution of gold responses often diverges significantly from the model’s own response distribution. For instance, consider an example from the MBPP dataset in Table 1. The base model, Mistral-7B-Instruct-v0.2, assigns a log probability of $-84.32$ to the gold answer. When prompted with the same question, the model generates a response containing the same information as the gold output, but with a much higher log probability of $-36.64$. This phenomenon is common in generation tasks, where semantically equivalent responses can have widely varying log-likelihood scores. This observation suggests that model-generated responses can align more closely with the model’s native distribution, whereas gold responses may lie further apart. As a result, training exclusively on gold responses risks introducing distributional drift, potentially reducing the model’s ability to generalize effectively.
@@ 2 Proposed Method
Let $\mathcal{M}_{\theta}$ parameterized by $\theta$ be a given large language model. Let $\theta=\theta_{0}$ be the given model weights obtained after pre-training and instruction tuning the model. We refer to $\mathcal{M}_{\theta_{0}}$ as the base model. Further, let $\mathcal{T}$ be the new task that we wish to teach the model $\mathcal{M}_{\theta}$, and let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ be the corresponding training dataset for $\mathcal{T}$. In standard SFT, we backpropagate through the standard Cross Entropy loss over the training dataset, $\mathcal{D}$. However, SFT can cause a degradation of $\mathcal{M}_{\theta_{0}}$ general capabilities by forcing the model to predict a gold response which is further away from the $\mathcal{M_{\theta}}$ responses’ distribution.
@@ 4 Results and Discussion
Dataset
Model
responses
Gold
paraphrases
Gold
responses
GSM8K 49.5% 42.1% 8.4%
MBPP 30.2% 32.4% 37.4%
NQ 66.0% 25.6% 8.1%
@@
Our findings from Tables 2 and 4 indicate that using the model’s response as the target rather than the gold response significantly enhances fine-tuning performance without compromising the model’s overall capabilities. Training on gold responses can cause a shift from the original distribution, negatively impacting the model’s generalization. Figure 2 illustrates this gap between the distributions of gold responses, gold paraphrases, and the base model’s responses. The distribution is plotted using $84$ random samples for which the model’s prediction and the paraphrases of the gold responses are acceptable. The plot shows that in several cases model responses can be valid and closer to model’s own distribution and therefore curating the training data in this way leads to minimal changes in the model parameters while fine-tuning. Thus, the closer the gold response’s paraphrases to the model distribution, the better it is (RQ3).
@@ A.1 Prompt for Generating the Response
Figure 3: Prompt used for training the model on GSM8K dataset.
We list the prompts used with mistral-instruct-v2 to generate the base model responses, gold paraphrases and training in this section. For the sake of consistency and fair comparison, the same prompts are used for fine-tuning using SFT, SDFT and S3FT techniques. Figure 4 and Figure 6 present the prompts used for generating the base model prediction and gold paraphrasing for MBPP dataset. To simplify the process of extracting the code in the generated output, we always ask the model to generate only the Python code starting with the string "[[BEGIN]]" and ending in the string "[[DONE]]". For the MBPP dataset, the training prompt is the same as the base model’s prediction prompt. Similarly, Figure 5 and Figure 7 present the prompt used for generating the base model response and paraphrasing the gold response. The details of the training prompt are provided in Figure 3.
@@ A.2 Datasets Details
GSM8K is a math word problem dataset comprising of 7473 training examples and 1319 test samples. Each question takes approximately 2-8 steps to solve. Solving these problems require knowledge of basic algebra only. Solutions are human written containing steps in natural language as well. We note that GSM8K does not have a validation set, so we took randomly sampled 150 samples from the dataset as validation.
@@ A.5 Judges for evaluating the model performance
For tasks like reading comprehension, conventional metrics like BLEU, ROUGE are useful but inadequate to capture the semantics of the generated responses. Therefore, we use LLM-as-a-judge for evaluation. Specifically, inspired by Badshah and Sajjad (2024), we use Mistral-instruct-v2 (7B) model as a judge with the prompt shown in Figure 8.

### v2

@@ 1 Introduction
SFT relies on gold responses for training. We make two key observations when performing SFT on LLMs: (1) in many datasets, model-generated responses—though differing from gold responses—are still valid and acceptable, and (2) the distribution of gold responses often diverges significantly from the model’s own response distribution. For instance, consider an example from the MBPP dataset in Table 1. The base model, Mistral-7B-Instruct-v0.2, assigns a log probability of $-84.32$ to the gold answer. When prompted with the same question, the model generates a response containing the same information as the gold output but with a much higher log probability of $-36.64$. This phenomenon is common in generation tasks, where semantically equivalent responses can have widely varying log-likelihood scores. This observation suggests that model-generated responses can align more closely with the model’s native distribution, whereas gold responses may lie further apart. As a result, training exclusively on gold responses risks introducing distributional drift, potentially reducing the model’s ability to generalize effectively.
@@ 2 Proposed Method
Let $\mathcal{M}_{\theta}$ parameterized by $\theta$ be a given large language model. Let $\theta=\theta_{0}$ be the given model weights obtained after pre-training and instruction tuning the model. We refer to $\mathcal{M}_{\theta_{0}}$ as the base model. Further, let $\mathcal{T}$ be the new task that we wish to teach the model $\mathcal{M}_{\theta}$, and let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ be the corresponding training dataset for $\mathcal{T}$. In standard SFT, we backpropagate through the standard Cross Entropy loss over the training dataset, $\mathcal{D}$. However, SFT can cause a degradation of $\mathcal{M}_{\theta_{0}}$ general capabilities by forcing the model to predict a gold response which is further away from the $\mathcal{M}_{\theta_{0}}$ responses’ distribution.
@@ 4 Results and Discussion
Our findings from Tables 2 and 3 indicate that using the model’s response as the target rather than the gold response significantly enhances fine-tuning performance without compromising the model’s overall capabilities. Training on gold responses can cause a shift from the original distribution, negatively impacting the model’s generalization. Figure 2 illustrates this gap between the distributions of gold responses, gold paraphrases, and the base model’s responses. The distribution is plotted using $84$ random samples for which the model’s prediction and the paraphrases of the gold responses are acceptable. The plot shows that in several cases model responses can be valid and closer to model’s own distribution and therefore curating the training data in this way leads to minimal changes in the model parameters while fine-tuning. Thus, the closer the gold response’s paraphrases to the model distribution, the better it is (RQ3).
@@
Dataset
Model
responses
Gold
paraphrases
Gold
responses
GSM8K 49.5% 42.1% 8.4%
MBPP 30.2% 32.4% 37.4%
NQ 66.0% 25.6% 8.1%
@@ A.1 Prompt for Generating the Response
In this section, we list the prompts used with mistral-instruct-v2 to generate the base model responses, gold paraphrases, and training. For the sake of consistency and fair comparison, the same prompts are used for fine-tuning using SFT, SDFT and S3FT techniques. Figure 10 and Figure 6 present the prompts used for generating the base model prediction and gold paraphrasing for MBPP dataset. To simplify the process of extracting the code in the generated output, we always ask the model to generate only the Python code starting with the string "[[BEGIN]]" and ending in the string "[[DONE]]". For the MBPP dataset, the training prompt is the same as the base model’s prediction prompt. Figure 4 and Figure 7 present the prompt used for generating the base model response and paraphrasing the gold response. The details of the training prompt are provided in Figure 3. Similarly, Figures 5 and 9 show the prompts used for generating the base model output and gold paraphrasing, respectively.
@@ A.2 Datasets Details
GSM8K is a math word problem dataset comprising of 7473 training examples and 1319 test samples. Each question takes approximately 2-8 steps to solve. Solving these problem require knowledge of basic algebra only. Solutions are human written containing steps in natural language as well. We note that GSM8K does not have a validation set, so we took randomly sampled 150 samples from the dataset as validation.
@@
Figure 3: Prompt used for training the model on GSM8K dataset.
Figure 5: Prompt used for predicting the base model’s output on NQ dataset.
Figure 9: Prompt used for paraphrasing the gold response of training partition of the NQ dataset.
@@ A.5 Judges for evaluating the model performance
For tasks like reading comprehension, conventional metrics like BLEU and ROUGE are helpful but inadequate to capture the semantics of the generated responses. Thus, we employ a Language Model (LLM) as an evaluator. Specifically, inspired by Badshah and Sajjad (2024), we utilize Mistral-instruct-v2 (7B) model as a judge guided by the prompt illustrated in Figure 8.